OpenStudy (anonymous):

How much would $125 invested at 8% interest compounded annually be worth after 14 years? Round your answer to the nearest cent. Do not include units in your answer.

mathslover (mathslover):

so first of all : \( \begin{array}{} \cal{\large{\text{Hey! Good to see you on this site!}}}\\ \\ {\huge \mathtt{\color{blue}{\text{Welcome To OpenStudy!}}}} \\ \text{enjoy answering and asking questions here}\\ \cal{\large{\text{Tips:)}}} \\ \textbf{Ambassadors (people with A next to their name) are here to help you.}\\ \textbf{If you have any questions about the site, feel free to ask us!}\\ \textbf{Users who appear in Purple are Moderators and help run the site.}\\ \textbf{If you have a serious issue with the site or with the behavior of other}\\ \textbf{users}\\ \textbf{please feel free to raise it with them.}\\ \textbf{Want to study a particular subject? Click 'Find More Subjects'}\\ \textbf{search for it there.}\\ \textbf{If you get a good response from a user , click on best response}\\ \end{array}\) please also visit our openstudy coc : http://openstudy.com/code-of-conduct

mathslover (mathslover):

How much would $125 invested at 8% interest compounded annually be worth after 14 years? Do u know the formula for compound interest?

OpenStudy (callisto):

\[A = P (1+r)^n\] Where A is the amount, P is the principle, r is the interest rate, n is the number of period invested. So, put P = 125, r=8%(that is 0.08) and n = 14 into the formula. \[A = 125 (1+0.08)^{14} = ...?\]